OpenStudy (maheshmeghwal9):

200ml of 1M HCL is mixed with 300ml of 6M HCL & the final solution is diluted to 1000ml. Calculate molar concentration of \[[H^+]\] ion.

OpenStudy (foolaroundmath):

1M HCl => 1 L of solution contains 1 mole of HCl i.e. 1 mole of H+ ions. moles of H+ ions in 200ml of 1M HCl = 0.2 (1L contains 1 mole so 200 ml contains ? ) moles of H+ ions in 300ml of 6M HCl = 0.3*6 (6M => 6 moles of ions in 1 L solution) Total moles of H+ ions = 0.2+0.18 = 0.38 Final volume of solution = 1L So, final concentration of H+ = 0.38/1 = 0.38M

OpenStudy (maheshmeghwal9):

but @FoolAroundMath answer is 2M.

OpenStudy (anonymous):

i got your answer: V1=0,2 dm3 c1=1M V2=0,3 dm3 c2=6M V3= V1+V2 = 0,5 dm3 c1V1+c2V2=c3V3 c3= (c1V1+c2V2)/V3 c3= 1*0,2 + 6*0,3 /0,5 c3= 4M now lets make it simple and say c3 is new c1 c1= 4M V1= 0,5 dm3 V2=1 dm3 c1V1=c2V2 c2= c1V1/V2 c2= 4*0,5 / 1 = 2M

OpenStudy (maheshmeghwal9):

So @Kryten u mean that final concentration of solution is same as molar concentration of H+ ions ??????????????????

OpenStudy (anonymous):

because HCl is strong acid we take that it completly dissociates in water/solution and yes [H+] is equal to concentration of HCL.

OpenStudy (maheshmeghwal9):

oh thanx a lot:)

OpenStudy (foolaroundmath):

oh yeah .. my bad 3*0.6 = 0.18 ... good job me . rofl and yes it is 2M

OpenStudy (foolaroundmath):

3*0.6 = 1.8 total moles of H+ = 0.2+1.8 = 2 => concentration = 2/1 = 2M

OpenStudy (maheshmeghwal9):

oh u r correct too :) I only observed ur answer sorry.